Random variable X has a distribution:

P(x=0)=0.2 ; P(x=1)=0.3 ; P(x=2)=0.1 & P(x=3)=0.3 ; P(x=4)=0.1.

Find: a) E(x) and Var(x)

b) Find Fx(Xo)

c) Find quantile of order 1/4 and median

d) Find P(2<=x<=4)
Answer #1

given that

P(x=0)=0.2 ; P(x=1)=0.3 ; P(x=2)=0.1 & P(x=3)=0.3 ; P(x=4)=0.1

a)

E(X)=0.2*0+1*0.3+2*0.1+3*0.3+4*0.1

=1.8

E(X2)=0.2*02+12*0.3+22*0.1+32*0.3+42*0.1

=5

Now

Var(X)=E(X2) -[E(X)]2 =5-1.82=1.76

b)

F(x)=P(X\leq x)

now

F(x)=\left\{\begin{matrix} 0 &x<0 \\ 0.2 & x\leq 0\\ 0.5 & 0<x\leq 1\\ 0.6 & 1<x\leq 2 \\ 0.9 & 2<x\leq 3 \\ 1 & 3<x\leq 4 \end{matrix}\right.

c)

quantile of order 1/4th is 1st quartile that

is P(X<x) =0.25

since

P(X=0) =0.2

so it will take values between 0 and 1

here

P(X<1)=0.5

Hence 1 is median

d)

P(2\leq X\leq 4)=P(X=2)+P(X=3)+P(X=4)=0.1+0.3+0.1=0.5